A coach decides to choose 3 out of 5 athletes to compete in a race. How many different ways can the coach choose the athletes? Identify if the situation involves combination or permutations.
A 20; permutations
B 20; combinations
C 10; combinations
D 10; permutations

Answer by Theo(13342)

formula for number of combinations of athletes is:
n!/((n-x)!x!)
he uses 3 out of 5 athletes.
n = 5
x = 3
formula becomes:
5!/((5-3)!*3!) = (5*4*3*2*1)/((2*1)*(3*2*1)) = 10
-----
the possible combinations are:
-----
abc
abd
abe
acd
ace
ade
bcd
bce
bde
cde
-----